OpenStudy (anonymous):

* I'll answer your questions pertaining to English, Psych, and Philosophy if you help me with mine? What is the equation of the line passing through (-6, 1) and perpendicular to the line y = -3x + 1 in slope-intercept form? a.) y = 3x + 4 b.) y= 1/3 x +3 c.) y = -3x - 6 d.) y= -1/3 x - 5

OpenStudy (anonymous):

perpendicular to a line with slope \(-3\) means slope will be \(\frac{1}{3}\) the "negative reciprocal" there is only one choice of answer with slope \(\frac{1}{3}\)

OpenStudy (anonymous):

i.e. it is B no need for further computation

OpenStudy (anonymous):

Wait how did the negative reciprocal come to mind there?

OpenStudy (anonymous):

Viewing the problem I didn't even come across that thought

OpenStudy (anonymous):

I got b without the negative r

OpenStudy (anonymous):

Math is so complicated, how do you know this so well.....?

OpenStudy (anonymous):

because i know that if a line with slope \(m_1\) is perpendicular to a line with slope \(m_2\) then \[m_1m_2=-1\] i.e. \[m_1=-\frac{1}{m_2}\] and \[m_2=-\frac{1}{m_1}\]

OpenStudy (anonymous):

OHHHHHH it's asking for perpendicular, so the opposite....yes?

OpenStudy (anonymous):

Now i understand

OpenStudy (anonymous):

you have to know this to do the question. to find the equation of a line you need a slope and a point if you are not given the slope, you have to find it somehow

OpenStudy (anonymous):

\[m_1m_2=-1\]Where did this come from @satellite73

OpenStudy (anonymous):

@wio I think he is referring to the negative reciprocal

OpenStudy (anonymous):

I know the negative reciprocal, but he used that to show the negative reciprocal, not the other way around.

OpenStudy (anonymous):

You're right, I saw it as how it should've been

OpenStudy (anonymous):

Minor mistake on his part